OK, so this question is KILLING me:Jose made a $300 down payment on a used car. His monthly payment was $150, and he paid $3,900 in all. For how many months did he make payments?
i need what is called the Arithmetic Solution and the Algebratic solution.
let m be the number of months of payment.
i hope i explained this well enough. so um help? anyone?

Answer by jim_thompson5910(35256)

Arithmetic Solution:

First subtract off the down payment

3900-300 = 3600

So he's financed (or he borrowed) $3600

Then divide this by the monthly payment (since equal payments of $150 are made each month)

3600/150 = 24


So he made 24 payments.

This means that 24 months have elapsed.

----------------------------------------------------------------------
Algebraic solution:

(Down Payment) + (Monthly Payment)*(Number of Months) = Total Amount Paid

300 + 150*x = 3900

150*x = 3900-300

150*x = 3600

x = 3600/150

x = 24

So we get the same answer of 24

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------